OpenStudy (anonymous):

Use an appropriate method to evaluate each of the following definite integrals Lower limit: pi/4 Upper Limit: pi /2 Integral: xcos(3x)dx Use an appropriate method to evaluate each of the following definite integrals Lower limit: pi/4 Upper Limit: pi /2 Integral: xcos(3x)dx @Mathematics

OpenStudy (anonymous):

can we do integration by parts

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Please is you can show the steps I will greatly appreciate it, thank you kindly

OpenStudy (anonymous):

\[\int\limits_{\pi/4}^{\pi/2}xcos(x)dx = xsinx|_{\pi/4}^{\pi/2} - \int\limits_{\pi/4}^{\pi/2}\sin(x)dx = (xsin(x)+\cos(x))|_{\pi/4}^{\pi/2}\] Here I set u=x => du=dx and dv=cos(x)dx => v=sinx

OpenStudy (anonymous):

@kietjohn ok, so I am not suppose to end up with a numerical value?

myininaya (myininaya):

plug in upper limit then minus (plug in lower limit)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

So Kietjohn set up is correct?

myininaya (myininaya):

looks good to me

OpenStudy (anonymous):

K thanks

OpenStudy (anonymous):

I thought the dv should be cos(3x)

myininaya (myininaya):

yeah he did for cos(x) not cos(3x)